creative candles sells small candles for $3 each and large ones for $5 each. yesterday they collected $103. if they sold three more large cnadles than small ones, how many of each size candle were sold?

Answer by checkley79(3341)

3X FOR THE SMALL CANDLES.
5(X+3)
3X+5(X+3)=103
3X+5X+15=103
8X=103-15
8X=88
X=88/8
X=11 SMALL CANDLES WERE SOLD.
11+3=14 LARGE CANDLES WERE SOLD.
PROOF:
3*11+5(14)=103
33+70=103
103=103